OpenStudy (anonymous):

If f(x) = x + 4, find f(3). 1 7 −1 12

OpenStudy (anonymous):

7

OpenStudy (anonymous):

thank you... Could you explain how you figured it out? please?

OpenStudy (anonymous):

magic

OpenStudy (anonymous):

huehuehue... >:D

OpenStudy (anonymous):

i think thats how most people figure out math :P

OpenStudy (anonymous):

u got tht right

jimthompson5910 (jim_thompson5910):

f(x) = x + 4 f(3) = 3 + 4 .. replace every copy of x with 3 f(3) = 7

OpenStudy (anonymous):

@jim_thompson5910 did it right